# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 2.1 Markov Generator and Markov Semigroup
$\left[1+\lambda C_{b}\right]f(v)-\lambda a\cdot\partial_{v}f(v)-\lambda\mathscr{L}_{+}^{*}[f](v)=g(v)$ (2.16)
@@
(2.16) can be rewritten, for any $\tau\leq s$, as:
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\left[f(z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\right]-\lambda\mathscr{L}_{+}^{*}[f](z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}=g(z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}.$
Integrating along the characteristics, that is, integrating the previous equation in $s$ between $t_{0}\leq t$, we obtain:
$f(v)$ $=f(z_{v}(t_{0}-t))e^{-\int_{t_{0}}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}+\lambda\int_{t_{0}}^{t}\mathscr{L}_{+}^{*}[f](z_{v}(s-t))e^{-\int_{s}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\hskip 199.16928pt+\int_{t_{0}}^{t}g(z_{v}(s-t))e^{-\int_{s}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s.$
Being $t_{0}$ and $t$ arbitrary, we can choose $t=0$. We can also consider the limit $t\rightarrow-\infty$, which will make vanish the first term of the left hand side, using the fact that the exponential is smaller than $1$, $z_{v}(t_{0})=v-\lambda t_{0}a$ and $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Therefore, we obtain (at least formally):
$f(v)$ $=\lambda\int_{-\infty}^{0}\mathscr{L}_{+}^{*}[f](z_{v}(s))e^{-\int_{s}^{0}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s+\int_{-\infty}^{0}g(z_{v}(s))e^{-\int_{s}^{0}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$=\lambda\int_{0}^{+\infty}\mathscr{L}_{+}^{*}[f](z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s+\int_{0}^{+\infty}g(z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s,$
@@
$A[f]=\int_{0}^{+\infty}\mathscr{L}_{+}^{*}[f](z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$ (2.20)
@@
$B[g]=\int_{0}^{+\infty}g(z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s.$ (2.21)
@@
We first show that, for any $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $v\in\mathbb{R}^{d}$, $A[f](v)$ and $B[f](v)$ are well-defined quantities.
We have:
$\left|B[f]\right|(v)\leq\|f\|_{\infty}\int_{0}^{+\infty}e^{-s}\mathop{}\!\mathrm{d}s=\|f\|_{\infty},$ (2.22)
proving that $B[f]$ is a uniformly bounded function. As for $A$, we have:
$\left|A[f]\right|(v)$ $\leq\int_{0}^{+\infty}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|f\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq\frac{C_{b}\|f\|_{\infty}}{\lambda C_{b}}\int_{0}^{+\infty}[1+\lambda C_{b}]e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq-\frac{\|f\|_{\infty}}{\lambda}\left[e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\right]_{0}^{+\infty}=\frac{\|f\|_{\infty}}{\lambda},$ (2.23)
@@
Therefore, $A[f]$ and $B[f]$ are two continuous and bounded functions when $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$.
It remains to show that $A[f]$ and $B[f]$ are vanishing at infinity. In the case of $A$, for any $\varepsilon>0$, there exists $s_{0}=s_{0}(\lambda,\|f\|_{\infty},\varepsilon)>0$ large enough such that:
$\int_{s_{0}}^{+\infty}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|f\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)$ $e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq\frac{\|f\|_{\infty}}{\lambda}$ $e^{-\int_{0}^{s_{0}}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\leq\frac{\|f\|_{\infty}}{\lambda}e^{-s_{0}}\leq\varepsilon/2.$ (2.24)
In addition, since $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$, there exists $M\geq 0$ such that for all $v\in\mathbb{R}^{d}$ with $|v|\geq M$, we have for all $s\in[0,s_{0}]$ and $\omega\in\mathbb{S}^{d-1}$:
$|f((v+\lambda sa)^{\prime})|\leq\frac{\lambda\varepsilon}{2}\cdotp$ (2.25)
Therefore:
$\int_{0}^{s_{0}}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|f\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)$ $e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
@@
We have shown therefore that for $|v|$ large enough, we have $|A[f]|(v)\leq\varepsilon$, proving that $A[f]\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Similar arguments applied to the case of $B[f]$ provides the same conclusion for this operator, which concludes the proof of the lemma. ∎
In the end, we can consider a more general form of the equation $f-\lambda T[f]=g$, as summarized in the following result.
Proposition 2.4 (Generalized equation $f-\lambda T[f]=g$).
Let $\lambda>0$ be a strictly positive number, and let $g\in\mathcal{C}_{0}(\mathbb{R}^{d})$ be any continuous function vanishing at infinity. If $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ belongs to $\mathscr{D}(T)$ and solves the equation:
$f-\lambda T[f]=g,$ (2.27)
then $f$ solves the equation:
$f=\lambda A[f]+B[g],$ (2.28)
where the mappings $A$ and $B$ are defined in (2.20) and (2.21) respectively. Conversely, if $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ solves the equation (2.28) and if $f\in\mathscr{D}(T)$ or $g\in\mathscr{D}(T)$, then $f$ solves the equation (2.27).
@@
Finally, as a last intermediate step to establish the results we need to prove Theorem 1.1, we will prove that the operator $A$ is bounded, and determine its norm, as a linear operator acting on $\mathcal{C}_{0}(\mathbb{R}^{d})$.
@@
Let $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$. A similar estimate as in the proof of Lemma 2.3, slightly more precise, shows that, for any $v\in\mathbb{R}^{d}$:
$\left|A[f]\right|(v)$ $\leq C_{b}\|f\|_{\infty}\underbrace{\int_{0}^{+\infty}e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s}_{=I_{\lambda}(v)}.$ (2.29)
Observe that in the particular case $\gamma=0$, we have:
@@
$\left|\lambda A[f]\right|(v)\leq\frac{\lambda C_{b}}{1+\lambda C_{b}}\|f\|_{\infty},$ (2.31)
@@
In order to prove Theorem 1.1, we first establish that the closure $\overline{T}$ of the operator $T$, defined in (2.9), is a Markov generator. This will allow us to solve the adjoint equation (2.4) for the test functions $\varphi$ in the weak form (1.28) of (1.26).
@@
We will check all the assumptions of Definition 3 in the case of the unbounded operator T defined in (2.9).
@@
    The sequence of functions $(\varphi_{n})_{n}$ given by the expression (2.11) satisfy for $T$ the second condition of Definition 3, as a consequence of the different properties established in Lemma 2.2 for such functions. The second point of Definition 3 holds for $T$.
@@
    Let $\varphi\in\mathscr{D}(A)$ and let us first assume that $\inf_{v\in\mathbb{R}^{d}}\varphi(v)$ is not attained at infinity. Then there exists a compact set $K\subseteq\mathbb{R}^{d}$ such that
@@
    Then let $g=\varphi-\lambda A\varphi$, we have that:
    $\inf_{v\in\mathbb{R}^{d}}g(v)\leq g(v_{0})=\varphi(v_{0})-\lambda a\cdot\partial_{v}\varphi(v_{0})-\lambda\mathscr{L}^{*}(\varphi)(v_{0})\leq\varphi(v_{0})=\inf_{v\in\mathbb{R}^{d}}\varphi(v)$ (2.33)
@@
    Let $g=\varphi-\lambda A\varphi$, in this case for any $n\in\mathbb{N}^{*}$ we have:
    $\inf_{v\in\mathbb{R}^{d}}g(v)\leq g(v_{n})$ $=\varphi(v_{n})-\lambda a\cdot\partial_{v}\varphi(v_{n})-\lambda\mathscr{L}^{*}(\varphi)(v_{n})$
@@
    $\inf_{v\in\mathbb{R}^{d}}g(v)\leq\inf_{v\in\mathbb{R}^{d}}\varphi(v).$
@@
    As for the fourth point of Definition 3, the surjectivity of $\text{id}-\lambda\overline{T}$, where $\overline{T}$ is the closure of the operator $T$, is obtained in two steps. First, let us observe that the image of $\text{id}-\lambda T$ is dense in $\mathcal{C}_{0}(\mathbb{R}^{d})$. Indeed, let us consider any $g\in\mathscr{D}(T)$. By Lemma 2.3, the mapping $A$ introduced in (2.20) is a linear mapping from $\mathcal{C}_{0}(\mathbb{R}^{d})$ to itself and, by Proposition 2.5, $\text{id}-\lambda A$ is invertible. Hence we can find a unique $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ that solves the generalized equation (2.28). Since in addition $g$ was assumed to belong to $\mathscr{D}(T)$, by Proposition 2.4 the regularity of $g$ allows to deduce that $f\in\mathscr{D}(T)$ as well.
@@
To conclude, let us observe that $T$ being a Markov pregenerator, then so is its closure $\overline{T}$. Moreover $\overline{T}$ fulfills also the fourth condition of Definition 3 and we conclude that $\overline{T}$ is a Markov generator. Therefore, by the Hille-Yosida theorem 2.1, the operator $\bar{T}$ generates a Markov semigroup $S(t)$.
Furthermore if $\varphi_{0}\in\mathscr{D}(T)$ then $S(t)\varphi_{0}=\varphi(t,v)\in\mathcal{C}^{1}([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ can be chosen as a test function in Definition 1. Now for every $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ and $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ we define $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ by means of the duality formula
@@
Due to the existence of the semigroup $S(t)$ this definition is well-posed. Let now $u(t)\in\mathcal{C}^{1}([0,+\infty),\mathscr{D}(T))$. Notice that $\bar{T}u(t)=Tu(t)$ and that $S(t)(Tu(t))=T(S(t)u(t))$ since $u(t)\in\mathscr{D}(T)$ for all $t\geq 0$. Set
$\Phi(t,v)=\partial_{t}u(t,v)+Tu(t,v).$ (2.36)
@@
Integrating (2.37) with respect to $f_{0}$ over $\mathbb{R}^{d}$ and integrating over $[0,\bar{t}]\times\mathbb{R}^{d}$ yields
$\int_{0}^{\bar{t}}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)S(t)\Phi(t,v)=\int_{0}^{\bar{t}}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)\partial_{t}(S(t)u(t,v))=\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)(S(\bar{t})u(\bar{t},v)-u(0,v)).$ (2.38)
@@
$\int_{\mathbb{R}^{d}}f(\bar{t},\mathop{}\!\mathrm{d}v)u(\bar{t},v)$ $-\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)u(0,v)=\int_{0}^{\bar{t}}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)S(t)\Phi(t,v)=\int_{0}^{\bar{t}}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f(t,\mathop{}\!\mathrm{d}v)\Phi(t)$
$=\int_{0}^{\bar{t}}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f(t,\mathop{}\!\mathrm{d}v)(\partial_{t}u(t,v)+Tu(t,v))$ (2.39)
which proves that $f(t,\mathop{}\!\mathrm{d}v)$ is a weak solution.
@@
[…]
@@
The right-hand side is identically zero by (1.28), hence $g(t,v)=0$ for all $t$. We obtained therefore that $f_{1}=f_{2}$, hence the uniqueness.

### v2

@@ 2.1 Markov Generator and Markov Semigroup
$\left[1+\lambda C_{b}\right]\varphi(v)-\lambda a\cdot\partial_{v}\varphi(v)-\lambda\mathscr{L}_{+}^{*}[\varphi](v)=\psi(v)$ (2.16)
@@
(2.16) can be rewritten, for any $\tau\leq s$, as
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\left[\varphi(z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\right]-\lambda\mathscr{L}_{+}^{*}[\varphi](z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}=\psi(z_{v}(s))e^{\int_{\tau}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}.$
Integrating along the characteristics, that is, integrating the previous equation in $s$ between $t_{0}\leq t$, we obtain
$\varphi(v)$ $=\varphi(z_{v}(t_{0}-t))e^{-\int_{t_{0}}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}+\lambda\int_{t_{0}}^{t}\mathscr{L}_{+}^{*}[\varphi](z_{v}(s-t))e^{-\int_{s}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\hskip 199.16928pt+\int_{t_{0}}^{t}\psi(z_{v}(s-t))e^{-\int_{s}^{t}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s.$
The integration bounds $t_{0}$ and $t$ being arbitrary, we can choose $t=0$. We can in addition consider the limit $t_{0}\rightarrow$-$\infty$, which will make vanish the first term of the right hand side, using the fact that the exponential term is smaller than $1$, $z_{v}(t_{0})=v-\lambda t_{0}a$ and $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Therefore, we obtain (at least formally):
$\varphi(v)$ $=\lambda\int_{-\infty}^{0}\mathscr{L}_{+}^{*}[\varphi](z_{v}(s))e^{-\int_{s}^{0}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s+\int_{-\infty}^{0}\psi(z_{v}(s))e^{-\int_{s}^{0}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$=\lambda\int_{0}^{+\infty}\mathscr{L}_{+}^{*}[\varphi](z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s+\int_{0}^{+\infty}\psi(z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s,$
@@
$A[\varphi]=\int_{0}^{+\infty}\mathscr{L}_{+}^{*}[\varphi](z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$ (2.20)
@@
$B[\psi]=\int_{0}^{+\infty}\psi(z_{v}(-s))e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s.$ (2.21)
@@
We first show that, for any $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $v\in\mathbb{R}^{d}$, $A[\varphi](v)$ and $B[\varphi](v)$ are well-defined quantities.
We have
$\left|B[\varphi]\right|(v)\leq\|\varphi\|_{\infty}\int_{0}^{+\infty}e^{-s}\mathop{}\!\mathrm{d}s=\|\varphi\|_{\infty},$ (2.22)
proving that $B[\varphi]$ is a uniformly bounded function. As for $A$, we have
$\left|A[\varphi]\right|(v)$ $\leq\int_{0}^{+\infty}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|\varphi\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq\frac{C_{b}\|\varphi\|_{\infty}}{\lambda C_{b}}\int_{0}^{+\infty}[1+\lambda C_{b}]e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq-\frac{\|\varphi\|_{\infty}}{\lambda}\left[e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\right]_{0}^{+\infty}=\frac{\|\varphi\|_{\infty}}{\lambda},$ (2.23)
@@
Therefore, $A[\varphi]$ and $B[\varphi]$ are two continuous and bounded functions when $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$.
It remains to show that $A[\varphi]$ and $B[\varphi]$ are vanishing at infinity. In the case of $A$, for any $\varepsilon>0$, there exists $s_{0}=s_{0}(\lambda,\|\varphi\|_{\infty},\varepsilon)>0$ large enough such that
$\int_{s_{0}}^{+\infty}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|\varphi\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)$ $e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
$\leq\frac{\|\varphi\|_{\infty}}{\lambda}$ $e^{-\int_{0}^{s_{0}}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\leq\frac{\|\varphi\|_{\infty}}{\lambda}e^{-s_{0}}\leq\varepsilon/2.$ (2.24)
In addition, since $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$, there exists $M\geq 0$ such that for all $v\in\mathbb{R}^{d}$ with $|v|\geq M$, we have for all $s\in[0,s_{0}]$ and $\omega\in\mathbb{S}^{d-1}$:
$|\varphi((v+\lambda sa)^{\prime})|\leq\frac{\lambda\varepsilon}{2}\cdotp$ (2.25)
Therefore
$\int_{0}^{s_{0}}\left(\int_{\mathbb{S}^{d-1}}b\left(\frac{v+\lambda sa}{|v+\lambda sa|}\cdot\omega\right)|\varphi\left((v+\lambda sa)^{\prime}\right)|\mathop{}\!\mathrm{d}\omega\right)$ $e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s$
@@
We have shown therefore that for $|v|$ large enough, we have $|A[\varphi]|(v)\leq\varepsilon$, proving that $A[\varphi]\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Similar arguments applied to the case of $B[\varphi]$ provides the same conclusion for this operator, which concludes the proof of the lemma. ∎
In the end, we can consider a more general form of the equation $\varphi-\lambda T[\varphi]=\psi$, as summarized in the following result.
Proposition 2.4 (Generalized equation $\varphi-\lambda T[\varphi]=\psi$).
Let $\lambda>0$ be a strictly positive number, and let $\psi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ be any continuous function vanishing at infinity. If $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ belongs to $\mathscr{D}(T)$ and solves the equation:
$\varphi-\lambda T[\varphi]=\psi,$ (2.27)
then $\varphi$ solves the equation:
$\varphi=\lambda A[\varphi]+B[\psi],$ (2.28)
where the mappings $A$ and $B$ are defined in (2.20) and (2.21) respectively. Conversely, if $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ solves the equation (2.28) and if $\varphi\in\mathscr{D}(T)$ or $\psi\in\mathscr{D}(T)$, then $\varphi$ solves the equation (2.27).
@@
Finally, as a last intermediate step to establish the results we need for the proof of Theorem 1.1, we will prove that the operator $A$ is bounded, and determine its norm, as a linear operator acting on $\mathcal{C}_{0}(\mathbb{R}^{d})$.
@@
Let $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$. A similar estimate as in the proof of Lemma 2.3, slightly more precise, shows that, for any $v\in\mathbb{R}^{d}$:
$\left|A[\varphi]\right|(v)$ $\leq C_{b}\|\varphi\|_{\infty}\underbrace{\int_{0}^{+\infty}e^{-\int_{0}^{s}[1+\lambda C_{b}]\mathop{}\!\mathrm{d}u}\mathop{}\!\mathrm{d}s}_{=I_{\lambda}(v)}.$ (2.29)
We have:
@@
$\left|\lambda A[\varphi]\right|(v)\leq\frac{\lambda C_{b}}{1+\lambda C_{b}}\|\varphi\|_{\infty},$ (2.31)
@@
In order to prove Theorem 1.1, we first establish that the closure $\overline{T}$ of the operator $T$, defined in (2.9), is a Markov generator. This will allow us to solve the adjoint equation (2.4) for the test functions $\varphi$. In turn, choosing test functions $\varphi$ solving the adjoint equation (2.4) will yield a major simplification in the weak form (1.12) of the original equation (1.9).
@@
We will check the first three assumptions of Definition 3 in the case of the unbounded operator T defined in (2.9). Then, we will consider its closure $\overline{T}$, for which these assumptions will also hold, and prove finally that the fourth assumption of Definition 3 holds also for $\overline{T}$.
@@
    The sequence of functions $(\varphi_{n})_{n}$ given by the expression (2.11) satisfy for $T$ the second condition of Definition 3, as a consequence of the different properties established in Lemma 2.2 for such functions. Then, the second point of Definition 3 holds for $T$.
@@
    Let $\varphi\in\mathscr{D}(T)$ and let us first assume that $\inf_{v\in\mathbb{R}^{d}}\varphi(v)$ is not attained at infinity. It means that there exists a compact set $K\subseteq\mathbb{R}^{d}$ and $v_{0}\in K$ such that
@@
    Then let $\psi=\varphi-\lambda T[\varphi]$, we have that
    $\inf_{v\in\mathbb{R}^{d}}\psi(v)\leq\psi(v_{0})=\varphi(v_{0})-\lambda a\cdot\partial_{v}\varphi(v_{0})-\lambda\mathscr{L}^{*}(\varphi)(v_{0})\leq\varphi(v_{0})=\inf_{v\in\mathbb{R}^{d}}\varphi(v)$ (2.33)
@@
    Let $\psi=\varphi-\lambda T[\varphi]$, in this case for any $n\in\mathbb{N}^{*}$ we have
    $\inf_{v\in\mathbb{R}^{d}}\psi(v)\leq\psi(v_{n})$ $=\varphi(v_{n})-\lambda a\cdot\partial_{v}\varphi(v_{n})-\lambda\mathscr{L}^{*}(\varphi)(v_{n})$
@@
    $\inf_{v\in\mathbb{R}^{d}}\psi(v)\leq\inf_{v\in\mathbb{R}^{d}}\varphi(v).$
@@
    As for the fourth point of Definition 3, the surjectivity of $\text{id}-\lambda\overline{T}$, where $\overline{T}$ is the closure of the operator $T$, is obtained in two steps. First, let us observe that the image of $\text{id}-\lambda T$ is dense in $\mathcal{C}_{0}(\mathbb{R}^{d})$. Indeed, let us consider any $\psi\in\mathscr{D}(T)$. By Lemma 2.3, the mapping $A$ introduced in (2.20) is a linear mapping from $\mathcal{C}_{0}(\mathbb{R}^{d})$ to itself and, by Proposition 2.5, $\text{id}-\lambda A$ is invertible. Hence we can find a unique $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ that solves the generalized equation (2.28). Since in addition $\psi$ was assumed to belong to $\mathscr{D}(T)$, by Proposition 2.4 the regularity of $\psi$ allows to deduce that $\varphi\in\mathscr{D}(T)$ as well.
@@
To conclude, let us observe that $T$ being a Markov pregenerator, then so is its closure $\overline{T}$. Moreover $\overline{T}$ fulfills also the fourth condition of Definition 3 and we conclude that $\overline{T}$ is a Markov generator. Therefore, by the Hille-Yosida theorem 2.1, the operator $\overline{T}$ generates a Markov semigroup $S(t)$.
Furthermore if $\varphi_{0}\in\mathscr{D}(T)$ then $S(t)\varphi_{0}=\varphi(t,\cdot)\in\mathcal{C}^{1}$ can be chosen as a test function in Definition 1. Now for every $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ and $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ we define $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ by means of the duality formula
@@
Due to the existence of the semigroup $S(t)$ this definition of $f(t)$ is meaningful. Let now $u(t)\in\mathcal{C}^{1}([0,+\infty),\mathscr{D}(T))$. Notice that $\overline{T}[u(t)]=T[u(t)]$ and that $S(t)(T[u(t)])=T[S(t)u(t)]$ since $u(t)\in\mathscr{D}(T)$ for all $t\geq 0$. Set
$\Phi(t,v)=\partial_{t}u(t,v)+T[u(t)](v).$ (2.36)
@@
Integrating (2.37) with respect to $f_{0}$ over $\mathbb{R}^{d}$ and integrating over $0,t]\times\mathbb{R}^{d}$ yields
$\int_{0}^{t}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)S(s)\Phi(s,v)=\int_{0}^{t}\mathop{}\!\mathrm{d}s\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)\partial_{s}(S(s)u(s,v))=\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)(S(t)u(t,v)-u(0,v)).$ (2.38)
@@
$\int_{\mathbb{R}^{d}}f(t,\mathop{}\!\mathrm{d}v)u(t,v)$ $-\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)u(0,v)=\int_{0}^{t}\mathop{}\!\mathrm{d}s\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)S(s)\Phi(s,v)=\int_{0}^{t}\mathop{}\!\mathrm{d}t\int_{\mathbb{R}^{d}}f(s,\mathop{}\!\mathrm{d}v)\Phi(s)$
$=\int_{0}^{t}\mathop{}\!\mathrm{d}s\int_{\mathbb{R}^{d}}f(s,\mathop{}\!\mathrm{d}v)\big(\partial_{s}u(s,v)+T[u(s)](v)\big)$ (2.39)
which proves that $f(t,\mathop{}\!\mathrm{d}v)$ is a weak solution of equation (1.9), in the sense of Definition 1.
@@
$\begin{cases}-\frac{\mathop{}\!\mathrm{d}u}{\mathop{}\!\mathrm{d}t}(t)=a\cdot\partial_{v}u(t)+\mathscr{L}^{*}(u(t))&\\ u(\bar{t})=\varphi(\bar{t}),\;\;0\leq t\leq\bar{t}.\end{cases}$ (2.41)
@@
The right-hand side is identically zero by (1.12), hence $g(t)=0$ for all $t$. We obtained therefore that $f_{1}=f_{2}$, hence the uniqueness.